OpenStudy (anonymous):

solve: (x+2)(x+1)=X^2+11

OpenStudy (anonymous):

x^2+3x+2=x^2+11 ==>3x+2=11 ===> x=9/3=3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

ok got that right thanks

OpenStudy (anonymous):

yw